OpenStudy (anonymous):

rationalize the denominator \[14\sqrt{3}/3\sqrt{7}\]

OpenStudy (anonymous):

14 = 7x2 = sqrt7 x sqrt7 x 2

OpenStudy (anonymous):

14=7x2 =sqrt7 x ssqrt7 x 2

OpenStudy (anonymous):

\[10\sqrt{3} /3\sqrt{5}\]

OpenStudy (anonymous):

express 10 in terms of 5 and take it from there

OpenStudy (anonymous):

same thing as I explained above

OpenStudy (anonymous):

so is the answer\[2\sqrt{15}/3\]